Find the smallest value of parameter alpha such that equation

(sin(x))^2 * cos(2x) + alpha*(cos(x)^4-sin(x)^4)= -10(2* alpha +1)^2

has at least one real solution.


1
Expert's answer
2021-09-16T00:46:57-0400
"\cos^4x-\sin^4x=(\cos^2x-\sin^2x)(\cos^2x+\sin^2x)"

"=\cos(2x)\cdot1=\cos(2x)"

"\sin^2x=\dfrac{1-\cos(2x)}{2}"

"\dfrac{1-\cos(2x)}{2}(\cos(2x))+\alpha\cos(2x)=-10(2\alpha+1)^2"

"\cos^2(2x)-(1+2\alpha)\cos(2x)-20(2\alpha+1)^2=0"

"-1\leq \cos(2x)\leq 1"

"D=(-(1+2\alpha))^2-4(1)(-20(2\alpha+1)^2)=81(2\alpha+1)^2"

"\cos(2x)=\dfrac{(1+2\alpha)\pm9(1+2\alpha)}{2}"

"\cos(2x)=-4(1+2\alpha)\text{ or } \cos(2x)=5(1+2\alpha)"

"-1\leq-4(1+2\alpha)\leq1"

"-\dfrac{1}{4}\leq1+2\alpha\leq\dfrac{1}{4}"

"-\dfrac{5}{4}\leq2\alpha\leq -\dfrac{3}{4}"

"-\dfrac{5}{8}\leq\alpha\leq -\dfrac{3}{8}"


"-1\leq5(1+2\alpha)\leq1"

"-\dfrac{1}{5}\leq1+2\alpha\leq\dfrac{1}{5}"

"-\dfrac{6}{5}\leq2\alpha\leq -\dfrac{4}{5}"

"-\dfrac{3}{5}\leq\alpha\leq -\dfrac{2}{5}"

"-\dfrac{5}{8}=-\dfrac{25}{40}, -\dfrac{3}{5}=-\dfrac{24}{40}"



The smallest value of parameter "\alpha" is "-\dfrac{5}{8}" .